# Regret Analysis of Average-Reward Unichain MDPs via an Actor-Critic Approach

### v1

@@ 1.1 Related works
Actor-Critic Approaches: In contrast to direct policy gradient methods that estimate gradients from sampled trajectories, Actor-Critic (AC) methods use Temporal Difference (TD) based critics to aid policy gradient through bootstrapped value estimates. Further, direct methods scale poorly and rely on access to mixing or hitting times, typically unknown in practice. AC methods are more sample-efficient but introduce additional bias, making theoretical analysis harder. Regret results for AC methods are limited and focus on global convergence (pseudo-regret), as seen in [28, 36, 14]. The MLMC-NAC algorithm proposed in [14] achieves order-optimal convergence by leveraging multi-level Monte Carlo to reduce bias. However, both [14] and other analyses based on actor-critic methods assume ergodicity. Extending these results to unichain MDPs, which do not exhibit exponential mixing, requires fundamentally different techniques and forms the focus of this paper.
@@ 2 Problem Formulation and Preliminaries
A policy $\pi_{\theta}\in\Pi$ induces a transition function $P^{\pi_{\theta}}:\mathcal{S}\to\Delta^{|\mathcal{S}|}$, given as $P^{\pi_{\theta}}(s,s^{\prime})=\sum_{a\in\mathcal{A}}P(s^{\prime}|s,a)\pi_{\theta}(a|s)$, $\forall s,s^{\prime}\in\mathcal{S}$. The corresponding stationary distribution is defined as:
@@
where $\nu^{\pi_{\theta}}(s,a)\coloneqq d^{\pi_{\theta}}(s)\pi_{\theta}(a|s)$. This is commonly referred to as the Policy Gradient theorem and while the proof can be found in [32], we include it in Appendix C for completeness.
@@
where the symbol $\dagger$ denotes the Moore-Penrose pseudoinverse and $F(\theta)$ is the Fisher matrix as defined as
$F(\theta)=\operatorname{\mathbb{E}}\left[\nabla_{\theta}\log\pi_{\theta}(a|s)\otimes\nabla_{\theta}\log\pi_{\theta}(a|s)\right]$ (9)
where $\otimes$ symbolizes the outer product and the expectation is taken over ${(s,a)\sim\nu^{\pi_{\theta}}}$. Note that $\omega_{\theta}^{*}$ can be written as the minimizer of the function $L_{\nu^{\pi_{\theta}}}(\cdot,\theta)$ where $\forall\omega\in\mathbb{R}^{\mathrm{d}}$,
$L_{\nu^{\pi_{\theta}}}(\omega,\theta)\textstyle{=\dfrac{1}{2}\operatorname{\mathbb{E}}\left[\big{(}A^{\pi_{\theta}}(s,a)-\omega^{\top}\nabla_{\theta}\log\pi_{\theta}(a|s)\big{)}^{2}\right]}.$ (10)
@@ 3 Proposed Algorithm
We propose an Natural Actor-Critic algorithm with Batching (Algorithm 1), structured around $K$ outer iterations (or epochs). For clarity, we reserve the motivation for the update rules using exact policy gradients, value functions, and average rewards to Appendix B. Algorithm 1 uses noisy samples of these updates, which is described below.
@@
To approximate the value function $V^{\pi_{\theta}}(\cdot)$, we employ a linear critic $\hat{V}(\zeta,\cdot)=\zeta^{\top}\phi(\cdot)$, where $\zeta\in\mathbb{R}^{m}$ denotes the critic parameter and $\phi(s)\in\mathbb{R}^{m}$ is a feature mapping satisfying $|\phi(s)|\leq 1$ for all $s\in\mathcal{S}$. At each outer iteration $k$, we estimate $\xi_{k}=[\eta_{k},\zeta_{k}]^{\top}$, where $\eta_{k}$ approximates the average reward $J(\theta_{k})$ and $\zeta_{k}$ approximates the critic parameter $\zeta_{\theta_{k}}$.
The estimate $\xi_{k}$ is refined over $H$ inner-loop iterations. For each $h\in{0,\dots,H-1}$, we update
@@
where $\beta$ is the critic learning rate, $B$ is the batch size, and $z_{b}^{kh}=(s_{b}^{kh},a_{b}^{kh},s_{b+1}^{kh})$ is a sampled transition. The update direction $v(z_{b}^{kh};\xi_{h}^{k})$ follows standard constructions in Temporal Difference learning (e.g., [42]):
@@
and $c_{\beta}$ and $\beta$ denote the critic learning rates. The advantage estimate at sample $z_{b}^{kh}$ is computed as
@@
$b_{u}(\theta_{k},\xi_{k},z_{b}^{kh})=\hat{A}^{\pi_{\theta_{k}}}(z_{b}^{kh};\xi_{k})\nabla_{\theta}\log\pi_{\theta_{k}}(a_{b}^{kh}\mid s_{b}^{kh}),$ (16)
and refine the natural gradient estimate $\omega_{k}$ via the update:
@@
$A_{u}(\theta_{k},z_{b}^{kh})$ $=\nabla_{\theta}\log\pi_{\theta_{k}}(a_{b}^{kh}\mid s_{b}^{kh})\otimes\nabla_{\theta}\log\pi_{\theta_{k}}(a_{b}^{kh}\mid s_{b}^{kh}).$ (19)
After completing the inner updates, the policy is updated as: $\theta_{k+1}=\theta_{k}+\alpha\omega_{k},$ where $\alpha$ is the policy learning rate.
Notice that the algorithm operates using a single trajectory of total length $T=2KHB$. For each outer iteration $k$, the algorithm collects a contiguous segment of $2BH$ samples by following the current policy $\pi_{\theta_{k}}$, starting from the final state of the previous iteration. The first $BH$ samples from $\pi_{\theta_{k}}$ are used to perform $H$ iterations of the critic update subroutine, and the next $BH$ samples are used for $H$ iterations of the NPG update.
@@ 4.1 Assumptions
Let $M_{\theta}\coloneqq\mathbb{E}_{\theta}\left[\phi(s)\left(\phi(s)-\phi(s^{\prime})\right)^{\top}\right]$, where the expectation $\mathbb{E}_{\theta}$ is taken over $s\sim d^{\pi_{\theta}}$ and $s^{\prime}\sim P^{\pi_{\theta}}(s,\cdot)$. Then, for all $\theta$, there exists a constant $\lambda>0$ such that for all $x\in\mathrm{Ker}(M_{\theta})_{\perp}$ (i.e., the subspace orthogonal to the kernel of $M_{\theta}$), the following inequality holds:
@@ 4.2 Main Result
The choice of learning rates along with the regret bound containing all problem-specific constants is provided in Appendix I. We now provide a brief overview of the proof. We first begin with a regret decomposition obtained using standard techniques:
@@
The proof of the above is given in Appendix H. This regret decomposition crucially depends on the bias, $\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}\left[\omega_{k}\right]-\omega^{*}_{k})\|$, and second-order error of the NPG estimate, $\operatorname{\mathbb{E}}\|\omega_{k}-\omega_{k}^{*}\|^{2}$. These in turn dependent on the quality of the critic estimates. Hence the main focus of our proof consists of bounding these terms. The analysis of remaining terms then follows from these bounds; details are provided in Appendix I.
@@
Another consequence of Lemma 2 is in providing bounds for the value and $Q$ functions. In the ergodic case, it is known that the functions $V^{\pi}$ and $Q^{\pi}$ are $\operatorname*{\mathcal{O}}(t_{\mathrm{mix}})$. In the unichain case, using the above result, we show that these functions are still bounded, by $\operatorname*{\mathcal{O}}(C)$ instead. The proof of these results are given in Appendix D. We next describe the error analysis of the critic and NPG subroutines in Algorithm 1.
@@
$\mathrm{Ker}(A_{v}(\theta))\subseteq\mathrm{Ker}(A_{v}(z,\theta)).$ (25)
Equation (25) holds in the ergodic case, even if the critic matrix is singular, but generally does not hold in unichain settings (see Appendix F). However, it is satisfied when the Markov chain is restricted to its recurrent class. To address this, we show that, with high probability, the chain induced by $\pi_{\theta_{k}}$ enters the recurrent class, $\mathcal{S}^{\theta_{k}}_{R}$, very early, allowing us to focus on the recurrent class.
@@
Analysis under event $\mathcal{E}_{B}$: By standard Markov chain theory, $T_{\theta}$ is a stopping time hence, the Strong Markov Property (Theorem 1.2.5 of [25]) applies at $T_{\theta}$. In particular, conditioned on $s_{T}$, the post-$T_{\theta}$ trajectory is independent of the pre-$T_{\theta}$ path and evolves as a Markov chain initialized at $s_{T}$ under the original kernel $P^{\pi_{\theta}}$. As a result, it suffices to analyze the critic and NPG subroutines with the samples restricted to the irreducible class. In Appendix E, we analyze a generic stochastic linear recursion and show that these subroutines can be viewed as special cases of this recursion.
@@
$\|\Pi(\operatorname{\mathbb{E}}[\xi_{k}|\mathcal{E}_{B}]-\xi^{*}_{k})\|^{2}\leq\textstyle{\tilde{\operatorname*{\mathcal{O}}}\Bigg{(}e^{-c_{1}H}\|\xi_{0}-\xi^{*}\|^{2}+\frac{C_{\mathrm{tar}}^{2}m}{\lambda^{6}B^{2}}\Bigg{)}}$
where $c_{1}=\tfrac{\lambda^{3}}{16}$, $\xi_{k}^{*}=[A_{v}(\theta_{k})]^{\dagger}b_{v}(\theta_{k})=[\eta_{k}^{*},\zeta_{k}^{*}]^{\top}$ and $\Pi$ denotes the projection onto the space $\mathrm{Ker}(A_{v}(\theta))_{\perp}$.
The details regarding the verification of all conditions of the generic linear recursion used in Theorem 4 for the critic update is provided in Appendix F.
@@
$\textstyle{\|\operatorname{\mathbb{E}}[\omega_{k}|\mathcal{E}_{B}]-\omega_{k}^{*}\|^{2}\leq\operatorname*{\mathcal{O}}\Bigg{(}e^{-c_{2}H}\|\omega_{0}-\omega_{k}^{*}\|^{2}+\frac{dG_{1}^{8}C_{\mathrm{tar}}^{2}}{\mu^{6}B^{2}}+\frac{G_{1}^{2}\|\Pi(\operatorname{\mathbb{E}}[\xi_{k}|\mathcal{E}_{B}]-\xi_{k}^{*})\|^{2}}{\mu^{2}}+\frac{G_{1}^{2}\epsilon_{\mathrm{app}}}{\mu^{2}}\Bigg{)}}$
@@
Similar to the critic update, the details regarding the verification of all conditions of the generic linear recursion used in Theorem 4 for the NPG update is provided in Appendix G.

### v2

@@ 1.1 Related works
Actor-Critic Approaches: In contrast to direct policy gradient methods that estimate gradients from sampled trajectories, Actor-Critic (AC) methods use Temporal Difference (TD) based critics to aid policy gradient through bootstrapped value estimates. Further, direct methods scale poorly and rely on access to mixing or hitting times, typically unknown in practice. AC methods are more sample-efficient but introduce addtional bias, making theoretical analysis harder. Regret results for AC methods are limited and focus on global convergence (pseudo-regret), as seen in [29, 37, 16]. The MLMC-NAC algorithm in [16] achieves order-optimal convergence using multi-level Monte Carlo to reduce bias, but still assumes fast mixing. Extending such results to unichain MDPs, which lack exponential mixing, requires fundamentally different techniques and is the focus of this paper.
@@ 2 Problem Formulation and Preliminaries
A policy $\pi_{\theta}\in\Pi$ induces a transition function $P^{\pi_{\theta}}:\mathcal{S}\to\Delta^{|\mathcal{S}|}$, given as $P^{\pi_{\theta}}(s,s^{\prime})=\sum_{a\in\mathcal{A}}P(s^{\prime}|s,a)\pi_{\theta}(a|s),\quad\forall s,s^{\prime}\in\mathcal{S}$. The corresponding stationary distribution is defined as:
@@
where $\nu^{\pi_{\theta}}(s,a)\coloneqq d^{\pi_{\theta}}(s)\pi_{\theta}(a|s)$. This is commonly referred to as the Policy Gradient theorem and while the proof can be found in [33], we include it in Appendix B for completeness.
@@
where $\dagger$ denotes the Moore-Penrose pseudoinverse and $F(\theta)$ is the Fisher matrix defined as
$F(\theta)=\operatorname{\mathbb{E}}\left[\nabla_{\theta}\log\pi_{\theta}(a|s)\nabla_{\theta}\log\pi_{\theta}(a|s)^{\top}\right]$ (9)
where the expectation is taken over $(s,a)\sim\nu^{\pi_{\theta}}$}. This yields the NPG update
$\theta_{k+1}=\theta_{k}+\alpha\omega_{k}^{*}.$
@@ 3 Proposed Algorithm
We propose a Natural Actor–Critic with Batching algorithm (Algorithm 1), which runs for $K$ outer iterations (or epochs) of natural policy gradient updates. Each outer iteration includes $H$ inner iterations to estimate the average reward and value function, followed by another $H$ iterations to estimate the NPG direction. We first present the necessary preliminaries, followed by the detailed description of the algorithm.
Preliminaries: For any fixed $\theta_{k}$, it is known that $\omega_{k}^{*}$ is the solution to the optimization problem
$\omega_{k}^{*}=\arg\min_{\omega}L_{\nu^{\pi_{\theta_{k}}}}(\omega,\theta_{k}),$
where $L_{\nu^{\pi_{\theta}}}(\omega,\theta)$ is defined as follows
$L_{\nu^{\pi_{\theta}}}(\omega,\theta)$
$=\dfrac{1}{2}\operatorname{\mathbb{E}}\left[\big(A^{\pi_{\theta}}(s,a)-\omega^{\top}\nabla_{\theta}\log\pi_{\theta}(a|s)\big)^{2}\right],$ (11)
where the expectation is taken over $(s,a)\sim\nu^{\pi_{\theta}}$}. It can be seen that $L_{\nu^{\pi_{\theta}}}(\omega,\theta)$ is a quadratic function in $\omega$ with gradient given by $F(\theta)\omega-\nabla J(\theta)$. Since neither $F(\theta_{k})$ nor $\nabla J(\theta_{k})$ is available in closed form, we estimate them from samples and perform stochastic gradient descent on $L_{\nu^{\pi_{\theta_{k}}}}(\omega,\theta_{k})$ to obtain an approximate NPG direction $\omega_{k}$.
We estimate $F(\theta_{k})$ using the outer product $\nabla\log\pi_{\theta_{k}}(a|s)\nabla\log\pi_{\theta_{k}}(a|s)^{\top}]$ computed from samples $(s,a)\sim\nu^{\pi_{\theta_{k}}}$. However, estimating $\nabla J(\theta_{k})$ is more challenging, as it requires access to the advantage function, which is not directly observable. We note that $A^{\pi_{\theta_{k}}}(s,a)$ can be expressed using $J(\theta_{k})$ and $V^{\pi_{\theta_{k}}}$, as follows from the definition of $A^{\pi_{\theta_{k}}}(s,a)$ and Bellman’s equation. Consequently, we focus next on estimating these quantities.
Notice that $\eta_{k}=J(\theta_{k})$ is the solution to the following equation:
$R_{{\theta_{k}}}(\eta_{k})=0,\quad\text{where}\quad R_{{\theta_{k}}}(\eta)=\operatorname{\mathbb{E}}_{(s,a)\sim\nu^{\pi_{\theta_{k}}}}[\eta-r(s,a)].$ (12)
Given access to $R_{\theta_{k}}(\eta)$, the solution to the above can be computed via the following iterative update:
$\eta_{h+1}^{k}=\eta_{h}^{k}-c_{\beta}\beta R_{\theta_{k}}(\eta_{h}^{k}),$ (13)
where $\beta$ and $c_{\beta}$ are step-size parameters.
@@
To approximate the value function $V^{\pi_{\theta_{k}}}(\cdot)$, we employ a linear critic $\hat{V}(\zeta_{\theta_{k}},\cdot)=\zeta_{\theta_{k}}^{\top}\phi(\cdot)$, where $\zeta_{\theta_{k}}\in\mathbb{R}^{m}$ denotes the critic parameter and $\phi(s)\in\mathbb{R}^{m}$ is a feature mapping satisfying $|\phi(s)|\leq 1$ for all $s\in\mathcal{S}$. We also assume that the feature vectors $\phi(s)$} are linearly independent. The parameter $\zeta_{\theta_{k}}\in\mathbb{R}^{m}$ is obtained by solving the following optimization problem:
$\min_{\zeta\in\mathbb{R}^{m}}E(\theta_{k},\zeta)\coloneqq\frac{1}{2}\operatorname{\mathbb{E}}_{s\sim d^{\pi_{\theta_{k}}}}\left(V^{\pi_{\theta_{k}}}(s)-\hat{V}(\zeta,s)\right)^{2}.$ (14)
This formulation enables the use of a gradient-based iterative procedure to compute $\zeta_{\theta_{k}}$. Specifically, the gradient of $E(\theta_{k},\zeta)$ is given by:
$\begin{split}\nabla_{\zeta}E(\theta_{k},\zeta)&=\operatorname{\mathbb{E}}_{s\sim d^{\pi_{\theta_{k}}}}\left[\left(V^{\pi_{\theta_{k}}}(s)-\hat{V}(\zeta,s)\right)\nabla_{\zeta}\hat{V}(\zeta,s)\right]\\ &=\operatorname{\mathbb{E}}_{s\sim d^{\pi_{\theta_{k}}}}\left[\left(V^{\pi_{\theta_{k}}}(s)-\zeta^{\top}\phi(s)\right)\phi(s)\right]\end{split}$ (15)
The above quantity can be approximated as follows
$\nabla_{\zeta}E(\theta_{k},\zeta)$ $=\operatorname{\mathbb{E}}_{s\sim d^{\pi_{\theta_{k}}}}\left[\left(V^{\pi_{\theta_{k}}}(s)-\zeta^{\top}\phi(s)\right)\phi(s)\right]$ (16)
$=\operatorname{\mathbb{E}}_{(s,a)\sim\nu^{\pi_{\theta_{k}}},s^{\prime}\sim P(\cdot|s,a)}\left[\left(r(s,a)-J(\theta_{k})+V^{\pi_{\theta_{k}}}(s^{\prime})-\zeta^{\top}\phi(s)\right)\phi(s)\right]$ (17)
$\approx\operatorname{\mathbb{E}}_{(s,a)\sim\nu^{\pi_{\theta_{k}}},s^{\prime}\sim P(\cdot|s,a)}\left[\left(r(s,a)-J(\theta_{k})+\zeta^{\top}\phi(s^{\prime})-\zeta^{\top}\phi(s)\right)\phi(s)\right]$ (18)
$\coloneqq\widetilde{\nabla}_{\zeta}E(\theta_{k},\zeta)$ (19)
Using this gradient, $\zeta_{\theta_{k}}$ can be computed iteratively via:
$\zeta_{h+1}^{k}=\zeta_{h}^{k}-\beta\widetilde{\nabla}_{\zeta}E(\theta_{k},\zeta_{h}^{k}).$ (20)
The updates in (13), and (20) can be combined into one update for $\xi_{h}^{k}=[(\eta^{k}_{h})^{\top},(\zeta_{h}^{k})^{\top}]^{\top}$ as follows
$\xi_{h+1}^{k}=\xi_{h}^{k}-\beta[c_{\beta}R_{\theta}(\eta_{h}^{k})^{\top},\widetilde{\nabla}_{\zeta}E(\theta_{k},\zeta_{h}^{k})^{\top}]^{\top}$ (21)
The joint update rule for $\eta^{k}_{h}$ and $\zeta^{k}_{h}$ in (21) forms the basis of the critic step in Algorithm 1, where sample averages from the trajectory are used in place of the intractable expectations.
Algorithm 1 operates using a single trajectory of total length $T=2KHB$. For each outer iteration $k$, the algorithm collects a contiguous segment of $2BH$ samples by following the current policy $\pi_{\theta_{k}}$, starting from the final state of the previous iteration. The first $BH$ samples from $\pi_{\theta_{k}}$ are used to perform $H$ iterations of the critic update subroutine, and the next $BH$ samples are used for $H$ iterations of the NPG update.
Critic subroutine: For each $h\in{0,\dots,H-1}$, we use a sample averaged version of (21) to obtain the update
@@
where $\beta$ is the critic learning rate, $B$ is the batch size, $z_{b}^{kh}=(s_{b}^{kh},a_{b}^{kh},s_{b+1}^{kh})$ is a sampled transition, and $v(z_{b}^{kh};\xi_{h}^{k})$ is given by:
@@
The resulting update follows standard constructions in Temporal Difference learning (e.g., [44]). After $H$ critic updates, we take $\eta_{k}\coloneqq\eta_{H}^{k}$ and $\zeta_{k}\coloneqq\zeta_{H}^{k}$ as the final estimates of the average reward and critic parameters, giving $\hat{V}(\zeta_{k},\cdot)=\zeta_{k}^{\top}\phi(\cdot)$ as the value estimate.
Natural Policy Gradient estimation: The advantage estimate at sample $z_{b}^{kh}$ is computed from the Bellman equation using $\eta_{k}$ and $\hat{V}(\zeta_{k},\cdot)$ as follows
@@
$b_{u}(\theta_{k},\xi_{k},z_{b}^{kh})=\hat{A}^{\pi_{\theta_{k}}}(z_{b}^{kh};\xi_{k})\nabla_{\theta}\log\pi_{\theta_{k}}(a_{b}^{kh}|s_{b}^{kh}),$ (26)
and refine the natural gradient estimate $\omega_{k}$ via $H$ iterations of the update:
@@
$A_{u}(\theta_{k},z_{b}^{kh})$ $=\nabla_{\theta}\log\pi_{\theta_{k}}(a_{b}^{kh}|s_{b}^{kh})\nabla_{\theta}\log\pi_{\theta_{k}}(a_{b}^{kh}|s_{b}^{kh})^{\top}.$ (29)
Set $\omega_{k}\coloneqq\omega_{H}^{k}$. The policy is updated as: $\theta_{k+1}=\theta_{k}+\alpha\omega_{k},$ where $\alpha$ is the policy learning rate.
@@ 4.1 Assumptions
Let $M_{\theta}\coloneqq\mathbb{E}_{\theta}\left[\phi(s)\left(\phi(s)-\phi(s^{\prime})\right)^{\top}\right]$, where the expectation $\mathbb{E}_{\theta}$ is taken over $s\sim d^{\pi_{\theta}}$ and $s^{\prime}\sim P^{\pi_{\theta}}(s,\cdot)$. Then, for all $\theta$, there exists a constant $\lambda>0$ such that for all $x\in\ker(M_{\theta})^{\perp}$ (i.e., the subspace orthogonal to the kernel of $M_{\theta}$), the following inequality holds:
@@ 4.2 Main Result
The choice of learning rates along with the regret bound containing all problem-specific constants is provided in Appendix H. We now provide a brief overview of the proof. We first begin with a regret decomposition obtained using standard techniques:
@@
The proof of the above is given in Appendix G. This regret decomposition crucially depends on the bias, $\operatorname{\mathbb{E}}\|(\operatorname{\mathbb{E}}_{k}\left[\omega_{k}\right]-\omega^{*}_{k})\|$, and second-order error of the NPG estimate, $\operatorname{\mathbb{E}}\|\omega_{k}-\omega_{k}^{*}\|^{2}$. These in turn dependent on the quality of the critic estimates. Hence the main focus of our proof consists of bounding these terms. The analysis of remaining terms then follows from these bounds; details are provided in Appendix H.
@@
Another consequence of Lemma 2 is in providing bounds for the value and $Q$ functions. In the ergodic case, it is known that the functions $V^{\pi}$ and $Q^{\pi}$ are $\operatorname*{\mathcal{O}}(t_{\mathrm{mix}})$. In the unichain case, using the above result, we show that these functions are still bounded, by $\operatorname*{\mathcal{O}}(C)$ instead. The proof of these results are given in Appendix C. We next describe the error analysis of the critic and NPG subroutines in Algorithm 1.
@@
$\ker(A_{v}(\theta))\subseteq\ker(A_{v}(z,\theta)).$ (35)
Equation (35) holds in the ergodic case, even if the critic matrix is singular, but generally does not hold in unichain settings (see Appendix E). However, it is satisfied when the Markov chain is restricted to its recurrent class. To address this, we show that, with high probability, the chain induced by $\pi_{\theta_{k}}$ enters the recurrent class, $\mathcal{S}^{\theta_{k}}_{R}$, very early, allowing us to focus on the recurrent class.
@@
Analysis under event $\mathcal{E}_{B}$: By standard Markov chain theory, $T_{\theta}$ is a stopping time hence, the Strong Markov Property (Theorem 1.2.5 of [27]) applies at $T_{\theta}$. In particular, conditioned on $s_{T}$, the post-$T_{\theta}$ trajectory is independent of the pre-$T_{\theta}$ path and evolves as a Markov chain initialized at $s_{T}$ under the original kernel $P^{\pi_{\theta}}$. As a result, it suffices to analyze the critic and NPG subroutines with the samples restricted to the irreducible class. In Appendix D, we analyze a generic stochastic linear recursion and show that these subroutines can be viewed as special cases of this recursion.
@@
[…]
where $c_{1}=\tfrac{\lambda^{3}}{16}$, $\xi_{k}^{*}=[A_{v}(\theta_{k})]^{\dagger}b_{v}(\theta_{k})=[\eta_{k}^{*},\zeta_{k}^{*}]^{\top}$ and $\Pi$ denotes the projection onto the space $\ker(A_{v}(\theta))^{\perp}$.
The details regarding the verification of all conditions of the generic linear recursion used in Theorem 4 for the critic update is provided in Appendix E.
@@
[…]
@@
Similar to the critic update, the details regarding the verification of all conditions of the generic linear recursion used in Theorem 4 for the NPG update is provided in Appendix F.